The mean science achievement score of a group of Grade 8 students is 78 and s of 10. Granting that the scores of these students are normally distributed, how many would you expect to score less than 85? If the number of students who took the test is 200, how many would that be? How many students would score more than 85?
Answer by stanbon(48568)

The mean science achievement score of a group of Grade 8 students is 78 and
s of 10.
Granting that the scores of these students are normally distributed, how many would you expect to score less than 85?
Find the z-score of 85.
z(85)=(85-78)/10 = -0.7
P(0 < x < 85) = P(0 < z < -0.7) = 0.2580
-------------------------------------------------
If the number of students who took the test is 200, how many would that be? How many students would score more than 85?
-------
0.2580*200 = 51.6;rounding down you get 51 students scored below 85
200-51 = 149 students scored above 85
=================
Cheers,
Stan H.